OpenStudy (anonymous):

suppose the distance from the light source to the lens is 50 mm and the distance from the lens to the wall is 20 mm, what is the focal distance?

OpenStudy (anonymous):

Assuming your light source is your object, and your wall is your image, you can use the following equation: 1/f=1/do+1/di Or, the inverse of the focal distance (f) = the sum of the inverse of the distance of the object to the lens (do) plus the inverse of the distance of the image (di) to the lens. Hence (1/do+1/di)^-1=f (1/50mm + 1/20mm)^-1=14mm